# CloudCast – Total Cloud Cover Nowcasting with Machine Learning

### v1

@@ 4 Discussion
CloudCast is capable of generating forecasts up to five hours ahead, though the later forecast periods are affected by increasing blurring. The observed blurring in the forecasts is consistent with prior research, which has highlighted challenges in forecasting small-scale cloud features using convolutional networks. This blurring is likely a limitation of the model, driven by the inherent unpredictability of atmospheric conditions and cloud dynamics. Tweaking hyperparameters or adjusting the depth of the network may offer minor improvements, but more recent studies have shown that more complex architectures, such as transformers, perform better when sufficient training data is available. Increasing the horizontal resolution of CloudCast could potentially sharpen the forecasts, but at the cost of introducing errors growing over time, as small-scale atmospheric dynamics are harder to predict. Another key limitation is the resolution of the ground truth data, which, at 4 km, is insufficient to fully capture summer-time convection, such as cumulus clouds. Improving the accuracy and resolution of cloud cover observations would likely enhance the training and performance of data-driven models like CloudCast.
A typical observation extrapolator cannot dynamically create and dissipate clouds. While CloudCast demonstrates some ability to generate new clouds, as seen in the case studies, the formations are often too weak and exhibit a temporal lag. Extending the skilful forecast period beyond three hours would require a stronger tendency for cloud creation and dissipation, which would likely necessitate a deeper understanding of cloud dynamics and the inclusion of additional weather-related parameters in the initial conditions, such as wind or moisture content.

### v2

@@ 4 Summary and discussion
CloudCast is capable of generating forecasts up to five hours ahead, though the later forecast periods are affected by increasing blurring. The observed blurring in the forecasts is consistent with prior research, which has highlighted challenges in forecasting small-scale cloud features using convolutional networks. This blurring is likely a limitation of the model, driven by the inherent unpredictability of atmospheric conditions and cloud dynamics. Tweaking hyperparameters or adjusting the depth of the network may offer minor improvements, but more recent studies have shown that more complex architectures, such as transformers, perform better when sufficient training data are available. Increasing the horizontal resolution of CloudCast could potentially sharpen the forecasts, but at the cost of introducing errors growing over time, as small-scale atmospheric dynamics are harder to predict. Another key limitation is the resolution of the ground truth data, which, at 4 km, is insufficient to fully capture summer-time convection, such as cumulus clouds. Improving the accuracy and resolution of cloud cover observations would likely enhance the training and performance of data-driven models like CloudCast.
A typical observation extrapolator cannot dynamically create and dissipate clouds. While CloudCast demonstrates some ability to generate new clouds, as seen in the case studies, the formations are often too weak and exhibit a temporal lag. Extending the skillful forecast period beyond three hours would require a stronger tendency for cloud creation and dissipation, which would likely necessitate a deeper understanding of cloud dynamics and the inclusion of additional weather-related parameters in the initial conditions, such as wind or moisture content.
CloudCast has only been tested over Northern Europe, but in principle, the same neural network architecture could applied to other regions with similar meteorological conditions, provided that sufficient cloud cover training data is available. Its transferability across different geostationary satellites is not a concern, as CloudCast does not use direct multispectral observations but rather post-processed cloud cover data. However, key limitations exist in generalizing these results. First, the ground truth data in the current domain contains artifacts due to its location at the edge of the satellite’s detectable area. These distortions, present during training, may not apply in regions closer to the satellite nadir, potentially affecting model performance. Second, atmospheric conditions play a crucial role: in Northern Europe, the dominant flow is from the southwest, and the model’s ability to generalize may be reduced in regions with significantly different prevailing wind patterns and cloud dynamics.